If y = Ce-2x + Cze2x is the general solution of the differential equation,what is the corresponding differential equation.
Answer #1

The genral sal of the differential equation is given as ce-in + Gera comparsing O with Yzcemin teman mi -2 m₂ 2 which has roo(M+2) (m2) = 0 m²- am tqm-420 ao М° -4 -4 = replacing m by differential ofrohor m=D=fn -D2-4y = 0 D²y-uyzo ry uy=0 da2 is So